A bullet loses 1/20 of its velocity after penetrating into a plank. How many planks are required to stop the bullet ?

Grade:11

4 Answers

Prudhvi teja
13 years ago

Dear mohammod

Let the thickness of one plank = d
and the acceleration provided by the plank = a

v^2 = vo^2 + 2ad
If n planks are required to stop the bullet, then
0^2 = vo^2 + 2a*nd
2and = -vo^2
n = vo^2/(-2ad) -----------------(1)

v = vo - vo/20 = 19 vo/20 in passing through one plank
(19 vo/20)^2 = vo^2 + 2ad
361/400 * vo^2 = vo^2 + 2ad
-2ad = vo^2(1 - 361/400)
-2ad = vo^2 * 39/400

Substituting this value of -2ad into equation (1):
n = vo^2/(vo^2 * 39/400) = 400/39
The minimum number of planks needed = smallest integer greater than 400/39 = 11
Ans: 11

 

Please feel free to post as many doubts on our discussion forum as you can.we will get you the answer and detailed  solution very  quickly.

All the best.

Regards,
Askiitians Experts
Prudhvi Teja

 

 


Now you can win exciting gifts by answering the questions on Discussion Forum. So help discuss any query on askiitians forum and become an Elite Expert League askiitian

vikas askiitian expert
13 years ago

applying v^2=u^2+2ad  ,here d is the width of plank

 v=u/20 after penetrating 0ne plank

  2ad=399u^2/400 .........eq1 

  for n number of planks total distance after which bullet stops is nd

 again applying v^2=u^2+2adn  and putting v=0

  n=u^2/2ad......eq2

 solving 1and 2 

   n=1

Susharitha
6 years ago
The formula is m=n^2÷2n-1Substituting in this we will be getting 400÷39So the answer should be approximately 10Therefore the required number of planks is 10....
Rishi Sharma
3 years ago
Dear Student,
Please find below the solution to your problem.

Let the thickness of one plank = d and the acceleration provided by the plank = a v^2 = vo^2 + 2ad
If n planks are required to stop the bullet,
then 0^2 = vo^2 + 2a*nd 2and = -vo^2 n = vo^2/(-2ad) -----------------(1)
v = vo - vo/20 = 19 vo/20
in passing through one plank
(19 vo/20)^2 = vo^2 + 2ad 361/400 * vo^2
= vo^2 + 2ad -2ad
= vo^2(1 - 361/400) -2ad
= vo^2 * 39/400
Substituting this value of -2ad into equation (1):
n = vo^2/(vo^2 * 39/400) = 400/39
The minimum number of planks needed = smallest integer greater than 400/39 = 11

Thanks and Regards